OpenStudy (anonymous):

Kevin's savings account balance changed from $1140 in January to $1450 in April. Find the average rate of change per month. Round your answer to the nearest dollar

jimthompson5910 (jim_thompson5910):

What's the change in dollars?

OpenStudy (anonymous):

please explain

jimthompson5910 (jim_thompson5910):

ie change in account balance

jimthompson5910 (jim_thompson5910):

you go from 1140 dollars to 1450 dollars so you added/subtracted how many dollars

OpenStudy (anonymous):

310

jimthompson5910 (jim_thompson5910):

so you added 310 dollars

jimthompson5910 (jim_thompson5910):

therefore, we know that the rate of change will be positive

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

january is the first month april is the 4th month so this change in balance occurred over 4-1 = 3 months

jimthompson5910 (jim_thompson5910):

so your rate of change is (change in balance)/(change in months) = 310/3 = ???

OpenStudy (anonymous):

okay thx i get it love

jimthompson5910 (jim_thompson5910):

yw